OpenStudy (anonymous):

Suppose you roll two number cubes and pick a letter of the alphabet at random. Find the probability you roll 2 even numbers and pick one of the vowels a, e, i, o, or u.

OpenStudy (triciaal):

incomplete independent events numbers and letters P(even cube 1 ) = 1/2 P(even cube 2) P(vowel) = 5/26 P(even on 2/even on 1) = 1/2*1/2 with vowel 1/4*5/26 = 5/104